# Stein Variational Evolution Strategies

### v1

@@ Abstract
Stein Variational Gradient Descent (SVGD) is a highly efficient method to sample from an unnormalized probability distribution. However, the SVGD update relies on gradients of the log-density, which may not always be available. Existing gradient-free versions of SVGD make use of simple Monte Carlo approximations or gradients from surrogate distributions, both with limitations. To improve gradient-free Stein variational inference, we combine SVGD steps with evolution strategy (ES) updates. Our results demonstrate that the resulting algorithm generates high-quality samples from unnormalized target densities without requiring gradient information. Compared to prior gradient-free SVGD methods, we find that the integration of the ES update in SVGD significantly improves the performance on multiple challenging benchmark problems.
@@ 1 Introduction
Figure 1: Illustration of Stein Variational ES. Multiple ES search distributions are updated in parallel based on the SVGD step.
@@
We consider the task of minimizing a function $f:{\mathbb{R}}^{d}\mapsto{\mathbb{R}}$. Since most real world problems – such as neural network parameter search – involve highly non-convex functions, the optimization process is very sensitive to its initialization (Sullivan et al., 2022; Li et al., 2018). To address this challenge, a common approach is to generate multiple candidate solutions and select the best one (Hernández-Lobato et al., 2014; Toussaint et al., 2024; Parker-Holder et al., 2020). This allows to cast the optimization task into an approximate inference problem which can be formulated as follows:
$\min_{q}D(q~{}\|~{}p)\qquad p(\mathbf{x})=\frac{e^{-f(\mathbf{x})}}{Z},$
@@
where the normalization constant $Z=\int_{{\mathbb{R}}^{d}}e^{-f(\mathbf{x})}\text{d}\mathbf{x}$ is typically untractable, $p$ and $q$ are probability distributions supported on ${\mathbb{R}}^{d}$, and $D$ is a suitable divergence, such as the Kullback-Leibler (KL) divergence.
Stein Variational Gradient Descent (SVGD) is a powerful method to approximate such a target distribution through a set of particles that is iteratively updated using gradient steps (Liu and Wang, 2016). This procedure makes SVGD a very sample efficient and flexible method that does not require long burn-in periods, nor makes assumptions about the family of the target distribution. Due to these properties, SVGD is a popular and powerful method for sampling and optimization (Lambert et al., 2021; Power and Berenson, 2024; Lee et al., 2023; Zhang et al., 2019).
@@
However, in many real-world problems – such as robotics and chemistry – the energy function $f$ is non-differentiable (Lambert et al., 2021; Englert and Toussaint, 2018; Maus et al., 2023). This non-differentiability poses a challenge for SVGD due to its reliance on the gradient of $\log p(\mathbf{x})\propto f(x)$. To facilitate Stein variational inference in these scenarios, prior work introduced a gradient-free version of SVGD that uses analytical gradients from a surrogate distribution that is fitted to the particle set at each step (Han and Liu, 2018, GF-SVGD). While this method comes with theoretical guarantees, fitting the surrogate to the objective function is challenging in practice (cf. Fig. 2). To circumvent these issues, other works instead used simple Monte Carlo (MC) gradients in the SVGD update (Liu et al., 2017; Lambert et al., 2021; Lee et al., 2023). However, the MC gradient estimate has high variance, which can lead to slow convergence on real-world problems.
@@
To address the aforementioned shortcomings of existing gradient-free SVGD methods, we propose a novel approach, Stein Variational Evolution Strategies (SV-ES) that bridges the domains of Evolution Strategies (ES) and distribution approximation. The idea of SV-ES is to replace the score term in the SVGD update by the search distribution mean update of an ES in order to improve gradient-free SVGD. We motivate our work based on prior results that established ES as a competitive alternative to gradient-based optimization algorithms, achieving higher performance and robustness on difficult objectives due to their inherent exploration capabilities (Salimans et al., 2017; Wierstra et al., 2014). In particular, the Covariance Matrix Adaptation Evolution Strategy (Hansen and Ostermeier, 2001, CMA-ES) is one of the most popular ES across many domains (Hansen and Kern, 2004; Hansen et al., 2010; Jankowski et al., 2023), due to its adaptive and efficient search process, which leverages natural gradient-like updates of its search distribution (Tian et al., 2021; Akimoto et al., 2012). Thus, we present Stein Variational CMA-ES (SV-CMA-ES) as a particular instance of SV-ES in this paper.
Our method, SV-CMA-ES, replaces the score term in the SVGD particle update by the search distribution mean update of CMA-ES. This means that our approach represents each SVGD particle by a separate ES search distribution, which is updated using the fitness values of all samples from the respective subpopulation and the estimated update steps of all other particles (cf. Figure 1). We validate the effectiveness of our proposed approach on challenging problems from multiple domains such as robot trajectory optimization and reinforcement learning. Our experimental results demonstrate that SV-CMA-ES improves considerably over prior gradient-free SVGD approaches. In particular, we show that our algorithm can be used to generate samples from challenging densities efficiently, as well as a blackbox optimizer of the underlying energy functions. We summarize our contributions as follows:
@@
    We introduce a new perspective on gradient-free SVGD by proposing a formulation that uses an ES to update the particles (Section 3.1).
@@
    We present SV-CMA-ES as a particular instance of our approach, which bypasses the need for a surrogate distribution of previous gradient-free SVGD approaches (Section 3.2).
@@
    We investigate the performance of our method on a wide range of problems and demonstrate that including ES in the gradient-free SVGD update significantly improves the optimization outcome (Section 5).
@@ 2.1 Stein Variational Gradient Descent
Stein Variational Gradient Descent (Liu and Wang, 2016, SVGD) is a non-parametric inference algorithm that approximates a target distribution with a set of particles $X=\{\mathbf{x}_{i}\}^{n}_{i=1}$ as $q(\mathbf{x})=\sum_{\mathbf{x}_{i}\in X}\delta(\mathbf{x}-\mathbf{x}_{i})/n$, where we use $\delta(\cdot)$ to denote the Dirac delta function. Given an initial set of particles, the goal is to determine an optimal particle transformation $\phi_{k}^{*}:{\mathbb{R}}^{d}\mapsto{\mathbb{R}}^{d}$ that maximally decreases the KL divergence $D_{\text{KL}}(q~{}\|~{}p)$:
$\begin{split}&\mathbf{x}_{i}\leftarrow\mathbf{x}_{i}+\epsilon\phi^{*}(\mathbf{x}_{i}),\quad\forall\mathbf{x}_{i}\in X\\ \text{s.t.\ }&\phi^{*}=\operatorname*{arg\,min}_{\phi\in{\cal F}}\left\{\left.\frac{d}{d\epsilon}D_{\text{KL}}(q_{[\epsilon\phi]}~{}\|~{}p)\right|_{\epsilon=0}\right\},\end{split}$ (1)
@@
The main result by Liu and Wang (2016) is the derivation of a closed form solution to this optimization problem. By choosing ${\cal F}$ as the unit sphere ${\cal B}_{k}$ of a vector-valued reproducing kernel Hilbert space ${\cal H}_{k}^{d}$, i.e.  ${\cal F}_{k}={\cal B}_{k}\{\phi\in{\cal H}_{k}^{d}:\lVert\phi\rVert_{{\cal H}_{k}^{d}}\leq 1\}$, with its kernel function $k(\cdot,~{}\cdot):{\mathbb{R}}^{d}\times{\mathbb{R}}^{d}\mapsto{\mathbb{R}}$, the authors show that the solution to Eq. (1) is:
$\phi^{*}_{k}(\cdot)\propto\operatorname*{\mathbb{E}}_{\mathbf{x}\sim q}\left[\nabla_{\mathbf{x}}\log p(\mathbf{x})k(\mathbf{x},\cdot)+\nabla_{\mathbf{x}}k(\mathbf{x},\cdot)\right].$ (2)
This result can be used to update the particle set iteratively using (1) and (2), where the expectation is estimated via MC approximation over the entire particle set $X$. Intuitively, the particle update balances likelihood maximization and particle repulsion. While the first term in Eq. (2) moves the particles towards regions of high likelihood, the latter counteracts this by repulsing particles based on the kernel function gradient, hence it has been called repulsive force (Ba et al., 2021; D’Angelo and Fortuin, 2021).
Because vanilla SVGD has been shown to be prone to the initialization of particles and mode collapse (Zhuo et al., 2018; Ba et al., 2021; Zhang et al., 2020), prior work proposed Annealed SVGD (Liu et al., 2017; D’Angelo and Fortuin, 2021). This extension of SVGD, reweighs the terms in the update based on the optimization progress (D’Angelo and Fortuin, 2021). Given the timestep-dependent annealing parameter $\gamma(t)\in{\mathbb{R}}$, the annealed update is:
@@ 2.2 Covariance Matrix Adaptation Evolution Strategy
The Covariance Matrix Adaptation Evolution Strategy (Hansen and Ostermeier, 2001, CMA-ES) is one of the most popular ES algorithms. We therefore choose it as starting point for our ES-based SVGD method. The core idea of the CMA-ES algorithm is to iteratively optimize the parameters of a Gaussian search distribution ${\cal N}(\mathbf{x},\sigma^{2}\mathbf{C})$ from which the solution candidates are sampled. While the algorithmic intuition of CMA-ES is similar to MC gradient approaches (Salimans et al., 2017), CMA-ES updates the search distribution following natural gradient steps, which has been shown to produce more efficient steps than standard gradient descent on multiple problems (Martens, 2020; Akimoto et al., 2012; Glasmachers et al., 2010).
Given a population of candidate samples $\boldsymbol{\xi}_{i}\sim{\cal N}(\mathbf{x},\sigma^{2}\mathbf{C})$, each iteration of CMA-ES updates the search parameters as follows. The mean of the search distribution is updated by recombining the current generation of candidates via a weighted sum:
$\mathbf{x}\leftarrow\mathbf{x}+\alpha_{\mathbf{x}}\sigma\textstyle\sum_{i=1}^{m}w_{i}\mathbf{y}_{i},\text{ where }\mathbf{y}_{i}=(\boldsymbol{\xi}_{i}-\mathbf{x})/\sigma.$ (4)
In this sum, the recombination weights $w_{i}$ are based on the rank of the fitness value $f(\boldsymbol{\xi}_{i})$ relative to all other candidates. For details on their exact computation, we refer to Hansen (2016). Next, the step-size $\sigma$ is updated based on the trace of prior steps. Given
$\hat{\mathbf{y}}=\textstyle\sum_{i=1}^{m}w_{i}\mathbf{y}_{i},~{}\lambda_{\text{eff}}=(\Sigma_{i=1}^{\lambda}w_{i}^{2})^{-1},\text{ and }$ (5)
$\mathbf{p}_{\sigma}\leftarrow(1-\alpha_{\sigma})\mathbf{p}_{\sigma}+\sqrt{\alpha_{\sigma}(2-\alpha_{\sigma})~{}\lambda_{\text{eff}}}~{}\mathbf{C}^{-\frac{1}{2}}\hat{\mathbf{y}},$ (6)
@@
$\sigma\leftarrow\sigma\times\exp\Bigl{(}\frac{\alpha_{\sigma}}{d_{\sigma}}\Bigl{(}\frac{\lVert\mathbf{p}_{\sigma}\rVert}{\operatorname*{\mathbb{E}}\lVert{\cal N}(0,{\rm\bf I})\rVert}-1\Bigr{)}\Bigr{)}.$ (7)
Here $\lambda\leq m$ is the elite population size hyperparameter that specifies how many candidates are considered in the update of the distribution parameters, and $d_{\sigma}$ is a smoothing parameter that controls the rate of step-size changes. This update guarantees that the step-size is increased iff the length of the optimization path $\mathbf{p}_{\sigma}$ exceeds what would be expected if steps were sampled randomly from a Gaussian. Finally, the covariance $\mathbf{C}$ is updated based on the covariance of the previous steps and current population fitness values:
@@
$\begin{split}&\mathbf{C}\leftarrow(1+d(h_{\sigma})\alpha_{1}-\alpha_{1}-\alpha_{\lambda}\textstyle\sum_{j=1}^{m}w_{j})\mathbf{C}\\ &\qquad+\alpha_{1}\mathbf{p}_{c}{\mathbf{p}_{c}}^{T}+\alpha_{\lambda}\textstyle\sum_{i=1}^{p}\bar{w}_{i}\mathbf{y}_{i}\mathbf{y}_{i}^{T},\end{split}$ (8)
@@
with $\mathbf{p}_{c}\leftarrow(1-\alpha_{c})\mathbf{p}_{c}+h_{\sigma}\sqrt{\alpha_{c}(2-\alpha_{c})\lambda_{\text{eff}}}~{}\hat{\mathbf{y}}_{i}.$ (9)
@@
In words, the covariance update performs exponential smoothing over the optimization path to update $\mathbf{C}$ based on the within- and between-step covariance of well performing solutions. Thus, Eq. (8) scales $\mathbf{C}$ along directions of successful steps, making the CMA-ES update follow a natural gradient direction to make the search converge faster (Akimoto et al., 2012). For completeness, we list the detailed update, including the definitions of $h_{\sigma},d(h_{\sigma})$ and $\bar{w}_{i}$ in the Appendix.
3 Stein Variational Evolution Strategies
This section introduces a novel framework of using a multi-population ES gradient-free SVGD. We first outline how ES can generally be used for gradient-free SVGD approximations. Then, we proceed by introducing the CMA-ES-based SVGD as a particular instance of this framework. The full algorithm is listed in Alg. 1. In the following, we use $n\in{\mathbb{N}}^{+}$ to refer to the number of particles that are used for SVGD inference, i.e. the number of ES search distributions that are run in parallel, each used sampling subpopulations of size $m\in{\mathbb{N}}^{+}$. This amounts to a total population size of $nm$ for ES-based algorithms.
3.1 SVGD via Evolution
Consider the task of sampling from a density $p(\mathbf{x})=e^{-f(\mathbf{x})}/Z$ with the untractable normalization constant $Z$. The SVGD update in Eq. (1) depends on $p$ solely via its score function:
$\nabla_{\mathbf{x}}\log p(\mathbf{x})=-\nabla_{\mathbf{x}}f(\mathbf{x}).$
We thus make the (trivial) observation that any algorithm, which estimates a gradient approximation $\tilde{\nabla}$, can be used to approximate the SVGD update (1) as follows:
$\mathbf{x}_{i}\leftarrow\mathbf{x}_{i}+\epsilon\operatorname*{\mathbb{E}}_{\mathbf{x}\sim q}\left[\tilde{\nabla}_{\mathbf{x}}f(\mathbf{x})k(\mathbf{x},\mathbf{x}_{i})+\nabla_{\mathbf{x}}k(\mathbf{x},\mathbf{x}_{i})\right].$ (10)
The idea of this work is thus to represent each SVGD particle by the mean of an ES search distribution and use the estimated steps of the ES algorithm to approximate the gradient in the SVGD particle update. Since ES are easily parallelizable on modern GPUs (Lange, 2023; Tang et al., 2022), this approach comes at small additional runtime given that objective function evaluations are not prohibitively expensive, as it is the case for many real world problems such as logistic regression or modern reinforcement learning simulators (Lange, 2022; Freeman et al., 2021).
One of the most straightforward instances of such a Stein Variational ES is based on the update that was proposed by Salimans et al. (2017). By parametrizing the particle $\mathbf{x}_{i}$ as a stochastic variable that is perturbed by Gaussian noise $\boldsymbol{\xi}_{i}\sim{\cal N}(0,\zeta^{2}{\rm\bf I})$, the stochastic finite difference gradient approximation that was proposed in the context of SVGD by Liu et al. (2017) is recovered:
~$\tilde{\nabla}_{\mathbf{x}_{i}}=\frac{1}{m}\sum_{j=1}^{m}f(\mathbf{x}_{i}+\boldsymbol{\xi}_{j}\zeta)\boldsymbol{\xi}_{j}/\zeta.$
3.2 Stein Variational CMA-ES

### v2

@@ Abstract
Efficient global optimization and sampling are fundamental challenges, particularly in fields such as robotics and reinforcement learning, where gradients may be unavailable or unreliable. In this context, jointly optimizing multiple solutions is a promising approach to avoid local optima. While Stein Variational Gradient Descent (SVGD) provides a powerful framework for sampling diverse solutions, its reliance on first-order information limits its applicability to differentiable objectives. Existing gradient-free SVGD variants often suffer from slow convergence and poor scalability. To improve gradient-free sampling and optimization, we propose Stein Variational CMA-ES, a novel gradient-free SVGD-like method that combines the efficiency of evolution strategies with SVGD-based repulsion forces. We perform an extensive empirical evaluation across several domains, which shows that the integration of the ES update in SVGD significantly improves the performance on multiple challenging benchmark problems. Our findings establish SV-CMA-ES as a scalable method for zero-order sampling and blackbox optimization, bridging the gap between SVGD and evolution strategies.
Stein Variational CMA-ES (SV-CMA-ES): Multiple ES Populations Optimize Fitness and Diversity
Quantitative Performance
Density Approximation
Reinforcement Learning
Qualitative Performance Solution Quality & Diversity
Refer to caption Refer to caption Ground Truth SV-CMA-ES Refer to caption Refer to caption Multi-Pop. CMA-ES (kernel-free) CMA-ES
Figure 1: Left: Illustration of Stein Variational CMA-ES. Multiple ES search distributions are updated in parallel, similar to the SVGD step. Middle: Quantitative comparison of different methods for sampling and RL control tasks. SV-CMA-ES obtains higher quality solutions than existing gradient-free SVGD-based approaches. Right: Qualitative comparison of different CMA-ES-based methods unveils that SV-CMA-ES generates more diverse samples than other methods. The full experimental details can be found in Section 5.
@@ 1 Introduction
Many optimization problems – such as neural network parameter search – involve highly non-convex objective functions, which makes the optimization process very sensitive to its initialization (Sullivan et al., 2022; Li et al., 2018). Thus, these hard optimization problems are commonly approached by generating multiple solution candidates from which the best is selected (Toussaint et al., 2024; Parker-Holder et al., 2020). This allows to frame the task of minimizing a function $f:{\mathbb{R}}^{d}\to{\mathbb{R}}$ as an approximate inference problem which can be formulated as follows:
$\min_{q}D(q\leavevmode\nobreak\ \|\leavevmode\nobreak\ p)\qquad p(\mathbf{x})=\frac{e^{-f(\mathbf{x})}}{Z},$
@@
where the normalization constant $Z=\int_{{\mathbb{R}}^{d}}e^{-f(\mathbf{x})}\text{d}\mathbf{x}$ is typically intractable, $p$ and $q$ are probability distributions supported on ${\mathbb{R}}^{d}$, and $D$ is a suitable divergence, such as the Kullback-Leibler (KL) divergence.
Stein Variational Gradient Descent (SVGD) is a powerful algorithm to solve this optimization problem through iteratively updating a particle set (Liu and Wang, 2016). As the approach is non-parametric and does not require the lengthy burn-in periods of Markov chain Monte Carlo (MCMC) methods (Andrieu et al., 2003), it is a computationally efficient method to approximate complex distributions. Due to these properties, SVGD is an increasingly popular first-order method for sampling and non-convex optimization (Zhang et al., 2019; Maken et al., 2021; Pavlasek et al., 2023).
@@
Unfortunately, the reliance of SVGD on the score function limits its applicability to differentiable objectives. In many real-world problems – such as robotics and chemistry – however, the energy function $f$ may not yield reliable gradients or be non-differentiable altogether (Lambert et al., 2021; Englert and Toussaint, 2018; Maus et al., 2023). To facilitate gradient-free Stein variational inference, prior work introduced a zero-order version of SVGD that uses analytical gradients from a surrogate distribution (Han and Liu, 2018, GF-SVGD). While the algorithm provably minimizes the KL divergence, fitting the surrogate to the objective function is challenging in practice, especially in higher dimensions (cf. Fig. 2). Alternatively, other works used simple Monte Carlo (MC) gradients in the SVGD update (Liu et al., 2017; Lambert et al., 2021; Lee et al., 2023). Again, this approach comes with limitations as the MC step estimate has high variance, which often leads to noisy updates and thus poor computational efficiency.
@@
To address the aforementioned shortcomings of existing gradient-free SVGD methods, we propose a novel approach, Stein Variational CMA-ES (SV-CMA-ES). Our method bridges the fields of Evolution Strategies (ES) and distribution approximation by updating multiple ES search distributions in parallel. The idea of SV-CMA-ES is to perform the distribution updates in a coordinated manner using a kernel-based repulsion term, which ensures an inter-population diversity similar to that in SVGD. We motivate our work based on prior results that established ES as a competitive alternative to gradient-based optimization algorithms, achieving higher performance and robustness on difficult objectives due to their inherent exploration capabilities (Salimans et al., 2017; Wierstra et al., 2014). In particular, the Covariance Matrix Adaptation Evolution Strategy (Hansen and Ostermeier, 2001, CMA-ES) is one of the most popular ES across many domains (Hansen et al., 2010; Jankowski et al., 2023), due to its adaptive and efficient search process, which leverages a dynamic step-size adaptation mechanism to increase convergence speeds (Akimoto et al., 2012).
We evaluate our proposed approach on a wide range of challenging problems from multiple domains, such as robot trajectory optimization and reinforcement learning. Our experimental results demonstrate that SV-CMA-ES improves considerably over existing gradient-free SVGD approaches. Fig. 1 summarizes our findings. Not only can our method be used to sample from challenging densities efficiently, but also as a blackbox optimizer on non-convex objectives. We outline our contributions as follows:
@@
    We introduce a novel zero-order method for diverse sampling and global optimization that combines ideas of SVGD with gradient-free ES, thus bypassing the need for a surrogate distribution required by previous gradient-free SVGD approaches (Section 3).
@@
    We validate our method, SV-CMA-ES, on a range of problems and demonstrate that it improves over prior gradient-free SVGD approaches in sampling and optimization tasks (Fig. 1 middle; Sec. 5.1-5.3).
@@
    We show that our presented method improves over prior CMA-ES-based methods because it combines the fast convergence rate of CMA-ES with the entropy-preserving optimization dynamics of SVGD (Fig. 1 right; Sec. 5.3).
@@ 2.1 Stein Variational Gradient Descent
Stein Variational Gradient Descent (Liu and Wang, 2016, SVGD) is a non-parametric inference algorithm that approximates a target distribution with a set of ϱ+italic-$\varrho\in{\mathbb{N}}^{+}$ particles $X=\{\mathbf{x}_{i}\}^{\varrho}_{i=1}$ as $q(\mathbf{x})=\sum_{\mathbf{x}_{i}\in X}\delta(\mathbf{x}-\mathbf{x}_{i})/\varrho$, where we use $\delta(\cdot)$ to denote the Dirac delta function. Given an initial set of particles, the goal is to determine an optimal particle transformation $\phi^{*}:{\mathbb{R}}^{d}\to{\mathbb{R}}^{d}$ that maximally decreases the KL divergence $D_{\text{KL}}(q\leavevmode\nobreak\ \|\leavevmode\nobreak\ p)$:
$\begin{split}&\mathbf{x}_{i}\leftarrow\mathbf{x}_{i}+\epsilon\phi^{*}(\mathbf{x}_{i}),\quad\forall\mathbf{x}_{i}\in X\\ \text{s.t.\ }&\phi^{*}=\operatorname*{arg\,min}_{\phi\in{\cal F}}\left\{\left.\frac{d}{d\epsilon}D_{\text{KL}}(q_{[\epsilon\phi]}\leavevmode\nobreak\ \|\leavevmode\nobreak\ p)\right|_{\epsilon=0}\right\},\end{split}$ (1)
@@
The main result by Liu and Wang (2016) is the derivation of a closed form solution to this optimization problem. By choosing ${\cal F}$ as the unit sphere ${\cal B}_{k}$ in a vector-valued reproducing kernel Hilbert space ${\cal H}_{k}^{d}$, i.e.  ${\cal F}_{k}={\cal B}_{k}\{\phi\in{\cal H}_{k}^{d}:\lVert\phi\rVert_{{\cal H}_{k}^{d}}\leq 1\}$, with its kernel function $k(\cdot,\cdot):{\mathbb{R}}^{d}\times{\mathbb{R}}^{d}\to{\mathbb{R}}$, the authors show that the solution to Eq. (1) is:
$\phi^{*}_{k}(\cdot)\propto\operatorname*{\mathbb{E}}_{\mathbf{x}\sim q}\bigl{[}\underbrace{\nabla_{\mathbf{x}}\log p(\mathbf{x})k(\mathbf{x},\cdot)}_{\text{driving force}}+\underbrace{\nabla_{\mathbf{x}}k(\mathbf{x},\cdot)}_{\text{repulsive force}}\bigr{]}.$ (2)
This result can be used to update the particle set iteratively using Eq. (1) and (2), where the expectation is estimated via MC approximation over the entire particle set $X$. Intuitively, the particle update balances likelihood maximization and particle repulsion: the first term drives particles toward regions of higher probability, while the second term counteracts this by repulsing particles based on the kernel gradient (D’Angelo and Fortuin, 2021; Ba et al., 2021).
Because vanilla SVGD is prone to the initialization of particles and mode collapse (Zhuo et al., 2018; Ba et al., 2021; Zhang et al., 2020), prior work proposed Annealed SVGD (Liu et al., 2017; D’Angelo and Fortuin, 2021). This extension of SVGD, reweighs the terms in the update based on the optimization progress (D’Angelo and Fortuin, 2021). Given the timestep-dependent temperature parameter $\gamma(t)\in{\mathbb{R}}$, the annealed update is:
@@ 2.2 Covariance Matrix Adaptation Evolution Strategy
The Covariance Matrix Adaptation Evolution Strategy (Hansen and Ostermeier, 2001, CMA-ES) is one of the most popular ES algorithms. We therefore choose it as the starting point for our ES-based SVGD method. The core idea of the CMA-ES algorithm is to iteratively optimize the parameters of a Gaussian search distribution ${\cal N}(\mathbf{x},\sigma^{2}\mathbf{C})$ from which the candidate solutions are sampled. While the algorithmic intuition of CMA-ES is similar to MC gradient approaches (Salimans et al., 2017), CMA-ES updates the search distribution following natural gradient steps, which has been shown to produce more efficient steps than standard gradient descent on multiple problems (Martens, 2020; Akimoto et al., 2012; Glasmachers et al., 2010).
We note that our notation in the following deviates from the default notation in the ES literature, as some of its variable names are typically associated with a different meaning compared to the variational inference (VI) literature. For instance, $\mu$ commonly is the symbol for the mean of a Gaussian in VI literature, while it refers to the number of elites in CMA-ES. To improve clarity, we thus use the variable $n$ to denote the size of a sampled CMA-ES population and $m$ for the number of selected elites. In our notation, CMA-ES is therefore an $(m,n)$ strategy. The CMA-ES algorithm relies on multiple hyperparameters which we fix to the default values from Hansen (2016). For completeness, we include the definitions of these variables – $w_{i},\alpha_{1},\alpha_{m},\alpha_{\sigma},h_{\sigma},d(h_{\sigma})$ and $\bar{w}_{i}$ – in the Appendix. Further, we slightly overload notation by using $\mathbf{p}$ to denote the evolution path updates following Hansen (2016) (unlike pdf’s which we denote by $p$).
Given a population of $n$ candidate samples $\boldsymbol{\xi}_{i}\sim{\cal N}(\mathbf{x},\sigma^{2}\mathbf{C})$, each iteration of CMA-ES updates the search parameters as follows. First, the samples $\boldsymbol{\xi}_{i}$ are evaluated and ranked by their fitness $f(\boldsymbol{\xi}_{i})$ in ascending order. To simplify notation, we assume ranked solutions in the following, i.e., we assume that index $i<j\rightarrow f(\boldsymbol{\xi}_{i})\leq f(\boldsymbol{\xi}_{j})$. This allows to assign each sample to a mutation weight $w_{i}$, where weights for better solutions are higher. For details on the exact computation of the weights, we refer to Hansen (2016) and to Appendix A of our work. The mean of the search distribution is then updated by mutating the $m\leq n$ best samples from the current generation of candidates:
$\mathbf{x}\leftarrow\mathbf{x}+\Delta\mathbf{x}_{\textsc{cma}}$ (4)
where
$\Delta\mathbf{x}_{\textsc{cma}}=\sigma\textstyle\sum_{i=1}^{m}w_{i}\mathbf{y}_{i},\text{ and }\mathbf{y}_{i}=(\boldsymbol{\xi}_{i}-\mathbf{x})/\sigma.$ (5)
Next, the parameters of the search distribution are updated. First, the step-size $\sigma$ is updated based on the history of prior steps. Given
$m_{\text{eff}}=(\Sigma_{i=1}^{m}w_{i}^{2})^{-1},\text{ and }$ (6)
$\mathbf{p}_{\sigma}\leftarrow(1-\alpha_{\sigma})\mathbf{p}_{\sigma}+\sqrt{\alpha_{\sigma}(2-\alpha_{\sigma})\leavevmode\nobreak\ m_{\text{eff}}}\leavevmode\nobreak\ \mathbf{C}^{-\frac{1}{2}}\Delta\mathbf{x}_{\textsc{cma}}/\sigma,$ (7)
@@
$\sigma\leftarrow\sigma\times\exp\Bigl{(}\frac{\alpha_{\sigma}}{d_{\sigma}}\Bigl{(}\frac{\lVert\mathbf{p}_{\sigma}\rVert}{\operatorname*{\mathbb{E}}\lVert{\cal N}(0,{\rm\bf I})\rVert}-1\Bigr{)}\Bigr{)},$ (8)
where $\mathbf{p}_{\sigma}$ is a moving average over the optimization steps, which unprojects the steps using $C^{-1/2}/\sigma$, so the resulting vector follows a standard normal. Thus, Eq. (8) automatically adapts the step-size based on the expected length of steps, similar to momentum-based optimizers (Kingma and Ba, 2015; Nesterov, 1983), with the hyperparameters $\alpha_{\sigma}$ and $d_{\sigma}$ governing the rate of the step-size changes. Finally, the covariance $\mathbf{C}$ is updated based on the covariance of the previous steps and current population fitness values:
@@
$\begin{split}&\mathbf{C}\leftarrow(1+\alpha_{1}d(h_{\sigma})-\alpha_{1}-\alpha_{m}\textstyle\sum_{j=1}^{n}w_{j})\mathbf{C}\\ &\qquad+\alpha_{1}\mathbf{p}_{c}{\mathbf{p}_{c}}^{T}+\alpha_{m}\textstyle\sum_{i=1}^{n}\bar{w}_{i}\mathbf{y}_{i}\mathbf{y}_{i}^{T},\end{split}$ (9)
with
@@
$\mathbf{p}_{c}\leftarrow(1-\alpha_{c})\mathbf{p}_{c}+h_{\sigma}\sqrt{\alpha_{c}(2-\alpha_{c})m_{\text{eff}}}\leavevmode\nobreak\ \Delta\mathbf{x}_{\textsc{cma}}/\sigma.$ (10)
@@
In words, the covariance update performs smoothing over the optimization path to update $\mathbf{C}$ based on the within- and between-step covariance of well-performing solutions. Thus, Eq. (9) scales $\mathbf{C}$ along directions of successful steps to make the search converge faster.
3 A Stein Variational Evolution Strategy
This section introduces a novel framework of using a multi-population ES for efficient discovery of multiple high-quality solutions to an optimization problem. The idea of this work is to represent each SVGD particle by the mean of an ES search distribution and use the estimated steps of the ES algorithm as the driving force in the SVGD particle update. Hence, our approach exploits the CMA-ES step-size adaptation mechanism to make gradient-free inference more efficient. Intuitively, the reformulated update permits larger particle updates, similar to momentum, especially in flat regions of the target. Since ES are easily parallelizable on modern GPUs (Lange, 2023; Tang et al., 2022), this approach comes at a small additional runtime cost. In the following, we use ϱ+italic-$\varrho\in{\mathbb{N}}^{+}$ to refer to the number of ES search distributions, $n\in{\mathbb{N}}^{+}$ to denote the size of each sampled population, and $m\in{\mathbb{N}}^{+}$ for the number of elite samples. This amounts to a total population size of $\varrho n$ for ES-based algorithms.